# Illuminating Darkness: Learning to Enhance Low-light Images In-the-Wild

### v2

@@ 1 Introduction
We propose TFFormer, a tuning-fork-shaped Transformer-CNN architecture that addresses key limitations in existing SLLIE methods. Unlike RGB-based [41, 6, 13] and Retinex-inspired [3, 50] models, TFFormer encodes luminance and chrominance separately to enable domain-specific feature learning for structure-preserving denoising and color-consistent enhancement. To bridge the inter-domain correlation gaps observed in recent fixed color-space methods [47], we introduce a Luminance–Chrominance Cross-Attention Block (LCCAB) and a joint decoder that enables adaptive fusion exclusively at the decoding stage. This late-stage interaction leverages spatially adaptive fusion, preserving domain-specific features while ensuring a structurally coherent reconstruction. Furthermore, we introduce the LC Guided Refinement Block (LCGRB) and LC Consistency Guidance loss (LCCG) to improve global structural fidelity and perceptual alignment. TFFormer outperforms the prior SOTA by +2.45 dB PSNR on LSD and generalizes well across diverse real-world scenes.
@@ 2.1 Scene acquisition
Align-pair Selection. For each scene, we matched the low-light input with two adjacent reference frames, generating two candidate pairs. We found that extreme low-light conditions significantly degraded the visibility of low-light inputs, making automatic alignment unreliable. To address this, we enhance the visibility of low-light images (for aligning purposes only) using histogram matching [31, 33], followed by an expert manual review to identify well-aligned pairs. Based on expert assessment, we discarded approximately 1,500 scenes exhibiting global misalignment. Fig.3(c) shows an example of DLI and NLI aligned-pairs.
@@ 2.2 Refining Training Pairs
Despite controlled references, minor spatial degradations like defocus, local darkness, or overexposed highlights may persist [17, 22]. These issues can occasionally cause deep models to produce blurry regions and inconsistent exposure in extreme cases (see Sec. 4.4). To improve training reliability, we introduce a reference-guided patch pair elimination method. Thus, we extracted $512\times 512$ patches, yielding 208,853 pairs, then filter-out  14% using two filters:
(1) Brightness threshold. We discarded patches whose reference images had average intensity below $T=10$, computed as
@@ 3.1 Luminance-Chrominance Mapping
where $\mathbf{F}_{L}$, $\mathbf{F}_{C}$, $\mathbf{I}_{B_{L}}$, $\mathbf{I}_{B_{C}}$in $\mathbb{R}^{H\times W\times 40}$. Retinex-based SLLIE models such as Retinexformer [3] compute boosted images directly from illumination features, biasing toward luminance and yielding colour over-enhancement. Our decoupled design retains chrominance information, leading to balanced enhancement with fewer colour artefacts.
@@ 3.5 LC Consistency Guidance (LCCG)
TFFormer exploits LC characteristics by computing a per-pixel $\ell_{1}$ distance between the predicted and ground-truth LC components to mitigate over-smoothing [32]. This distance is applied to both the intermediate reconstruction $I_{\text{rec}}$ and the final refinement $I_{\text{ref}}$. LCCG enforces consistency between the two stages, enhances color fidelity, sharpens structural details, and provides smoother gradients for stable training and progressive refinement.
@@ 4.1 Comparison between SLLIE methods
In addition to accuracy, TFFormer offers a favorable performance complexity trade-off. With just 5.87M parameters and 34.6 GFLOPs, it is significantly more efficient than recent methods like Diff-Retinex [50] (80.7M, 492.5 GFLOPs) over 13$\times$ larger and 14$\times$ more computational yet achieves markedly lower performance. TFFormer outperforms it by +3.56 dB in PSNR, +0.2005 in SSIM, and -0.1206 in LPIPS. Compared to the heaviest model, MIRNet [51] (31.79M, 816 GFLOPs), TFFormer is 5.4$\times$ smaller and 24$\times$ faster, while achieving +3.83 dB higher PSNR, +0.1088 higher SSIM, and -0.0417 lower LPIPS. These results underline TFFormer’s efficiency without compromising quality
@@ 4.3 Generalization across devices
Figure 8: Performance of LSD-TFFormer on diverse sources. Our method can enhance old social media images from eight years ago, even if they are heavily compressed. Best viewed in zoom.
@@ C.3 Training Details
We implemented our proposed TFFormer using PyTorch framework [29]. Our TFFormer optimized with an Adam optimizer [16] with a hyperparameter of $\beta_{1}=0.9$, $\beta_{2}=0.99$. We set the initial learning rate = $41e-$ 4 and adjusted it after training 100,000 steps using ReduceLROnPlateau scheduler [4]. We trained our model for 250,000 steps, utilizing a batch size of 12 and image dimensions $128\times 128\times 3$. Additionally, we leverage a low-light regularization coefficient $\lambda_{R}=0.2$ empharically to tune the proposed LC guidance. The training process spanned approximately 100 hours, executed on low-end hardware featuring an AMD Ryzen 3200G central processing unit (CPU) operating at 3.6 GHz, complemented by 32 GB of random-access memory, and an Nvidia GeForce GTX 3060 (12GB) graphical processing unit (GPU).

### v3

@@ 1 Introduction
We propose TFFormer, a tuning-fork-shaped Transformer-CNN architecture that addresses key limitations in existing SLLIE methods. Unlike RGB-based [41, 6, 13] and Retinex-inspired [3, 50] models, TFFormer leverages structural-chromatic information by encoding luminance and chrominance separately, enabling domain-specific feature learning for structure-preserving denoising and color-consistent enhancement. To bridge the inter-domain correlation gaps observed in recent fixed color-space methods [47], we introduce a Luminance–Chrominance Cross-Attention Block (LCCAB) and a joint decoder that enables adaptive fusion exclusively at the decoding stage. This late-stage interaction leverages spatially adaptive fusion, preserving domain-specific features while ensuring a structurally coherent reconstruction. Furthermore, we introduce the LC Guided Refinement Block (LCGRB) and LC Consistency Guidance loss (LCCG) to improve global structural fidelity and perceptual alignment. TFFormer outperforms the prior SOTA by +2.45 dB PSNR on LSD and generalizes well across diverse real-world scenes.
@@ 2.1 Scene acquisition
Align-pair Selection. Our LSD collection pipeline captures aligned image pairs by fixing the camera’s focal point in each scene. However, we observed that smartphone hardware occasionally shifts the focal point (10-35% of cases), requiring expert reviewers to evaluate spatial alignment and discard scenes with minor disparities. For each scene, we generated two candidate pairs by matching the low-light input with adjacent reference frames. To improve the visibility of extreme low-light inputs for evaluating spatial alignment, we applied histogram matching [31, 33] to boost visibility for manual expert review. Our experts discarded approximately 1,500 scenes with global misalignment. Fig.3(c) shows examples of aligned (selected) DLI and NLI pairs.
@@ 2.2 Refining Training Pairs
Despite controlled references, minor spatial degradations like defocus, local darkness, or overexposed highlights may persist [17, 22]. These issues can occasionally cause deep models to produce blurry regions and inconsistent exposure in extreme cases (see Sec. 4.4). Unlike existing works [42, 12] that ignore such critical spatial evaluation, we introduce a reference-guided patch pair elimination method. Thus, we extracted $512\times 512$ patches, yielding 208,853 pairs, then filter-out  14% using two filters: (1) Brightness threshold. We discarded patches whose reference images had average intensity below $T=10$, computed as
@@ 3.1 Luminance-Chrominance Mapping
where $\mathbf{F}_{L}$, $\mathbf{F}_{C}$, $\mathbf{I}_{B_{L}}$, $\mathbf{I}_{B_{C}}$ $\in$ $\mathbb{R}^{H\times W\times 40}$. Retinex-based SLLIE models such as Retinexformer [3] compute boosted images directly from illumination features, biasing toward luminance and yielding colour over-enhancement. Our decoupled design retains chrominance information, leading to balanced enhancement with fewer colour artefacts.
@@ 3.5 LC Consistency Guidance (LCCG)
TFFormer leverages LC characteristics by employing a per-pixel $\ell_{1}$ distance between predicted and ground-truth LC components to address over-smoothing [32]. Unlike traditional feature-level consistency losses, LCCG enables separate handling of structure and color information. This loss is applied to both the intermediate reconstruction $I_{\text{rec}}$ and the final refinement $I_{\text{ref}}$, ensuring consistency between stages. By enhancing color fidelity, sharpening structural details, and smoothing gradients, LCCG facilitates stable training and progressive refinement.
@@ 4.1 Comparison between SLLIE methods
[…]
@@ 4.3 Generalization across devices
Figure 8: Performance of LSD-TFFormer on diverse sources, including old, heavily compressed social media images.
@@ 5 Conclusion
6 Acknowledgment
This research was supported in part by the Alexander von Humboldt Foundation.
@@ C.3 Training Details
We implemented our proposed TFFormer using PyTorch framework [29]. Our TFFormer optimized with an Adam optimizer [16] with a hyperparameter of $\beta_{1}=0.9$, $\beta_{2}=0.99$. We set the initial learning rate = 1e41e-4 and adjusted it after training 100,000 steps using ReduceLROnPlateau scheduler [4]. We trained our model for 250,000 steps, utilizing a batch size of 12 and image dimensions $128\times 128\times 3$. Additionally, we leverage a low-light regularization coefficient $\lambda_{R}=0.2$ empharically to tune the proposed LC guidance. The training process spanned approximately 100 hours, executed on low-end hardware featuring an AMD Ryzen 3200G central processing unit (CPU) operating at 3.6 GHz, complemented by 32 GB of random-access memory, and an Nvidia GeForce GTX 3060 (12GB) graphical processing unit (GPU).